OpenStudy (chunkymonkay):

what is the y-intercept of the line with the equation 8x-1/3y=16 ([ ],0)

OpenStudy (chunkymonkay):

@dtan5457

OpenStudy (dtan5457):

Ok you need to put this in y=mx+b form

OpenStudy (dtan5457):

8x-1/3y=16 subtract the 8x -1/3y=16-8x

OpenStudy (dtan5457):

how do you make -1/3y to just y? you would multiply bu -3.

OpenStudy (chunkymonkay):

subtract the 8x from each side?

OpenStudy (dtan5457):

yeah, you need to isolate y

OpenStudy (dtan5457):

-1/3y=16-8x just multiply the whole thing by -3

OpenStudy (dtan5457):

then you will have it at y=mx+b form b (whole number) will be your intercept

OpenStudy (chunkymonkay):

okay give me a sec to see if I can do this

OpenStudy (chunkymonkay):

im just a little bit confused how would I multiply the whole thing by -3?

OpenStudy (dtan5457):

your probably looking it harder than it really is -1/3y=16-8x -1/3 times -3=y -8x times -3=24x etc

OpenStudy (chunkymonkay):

so the answer would be 24?

OpenStudy (dtan5457):

no. 24x is the slope, you still need to multiply 16 by -3

OpenStudy (chunkymonkay):

ohh okay

OpenStudy (dtan5457):

y=24x-?

OpenStudy (chunkymonkay):

48

OpenStudy (dtan5457):

well, -48

OpenStudy (chunkymonkay):

okay so the answer is -48

OpenStudy (dtan5457):

there you now have it at y=mx+b y=24x-48 -48 is your intercept

OpenStudy (dtan5457):

but if anything you have your coordinate placement wrong in your question

OpenStudy (dtan5457):

it should be 0,-48

OpenStudy (dtan5457):

-48,0 would define as a x axis intercept

OpenStudy (chunkymonkay):

that's how it looks on my test

OpenStudy (dtan5457):

well -48 is definitely the y intercept, not sure why your question words it like that

OpenStudy (chunkymonkay):

thank you

OpenStudy (dtan5457):

more questions?

OpenStudy (chunkymonkay):

yes I have 2 more

OpenStudy (dtan5457):

ask away